OpenStudy (anonymous):

ok i hav one i need to find the slope of the line i have the points on the line on the graph i cant dra it but the dots are on 0,6 and 7,-5

OpenStudy (anonymous):

i dont know ho to do wihtout seeing an equation

OpenStudy (aroub):

u will make an equation out of these points..

OpenStudy (aroub):

do u know how to find the slope?

OpenStudy (anonymous):

Slope when two points are given is (y2 -y1)/ (x2-x1) (0,6) (7,-5) slope = (6-(-5))/(0-7) = 11/(-7)= -11/7